For the real or imaginary nucleus of 74X175, find (a) the net electrical charge of the nucleus, (b) the number of neutrons, (c) the number of nucleons, (d) the approximate radius of the nucleus, and (e) the nuclear density.
Answer #1

(a)

Q=+74e

(b)

N=175-74= 101

(c)

N'=175

(d)

R=R_0A^{1/3}= 1.2\times (175)^{1/3}\: fm = 6.71\: fm

(e)

d=\frac {m}{V}=\frac {175\times 1.67\times 10^{-27}}{\frac {4}{3}\pi (6.71\times 10^{-15})^3}= 2.30\times 10^{17}\: kg/m^3